### v1

# AgMMU: A Comprehensive Agricultural Multimodal Understanding and Reasoning Benchmark

Abstract
We curate a dataset AgMMU for evaluating and developing vision-language models (VLMs) to produce factually accurate answers for knowledge-intensive expert domains. Our AgMMU concentrates on one of the most socially beneficial domains, agriculture, which requires connecting detailed visual observation with precise knowledge to diagnose, e.g., pest identification, management instructions, etc. As a core uniqueness of our dataset, all facts, questions, and answers are extracted from 116,231 conversations between real-world users and authorized agricultural experts. After a three-step dataset curation pipeline with GPT-4o, LLaMA models, and human verification, AgMMU features an evaluation set of 5,460 multiple-choice questions (MCQs) and open-ended questions (OEQs). We also provide a development set that contains 205,399 pieces of agricultural knowledge information, including disease identification, symptoms descriptions, management instructions, insect and pest identification, and species identification. As a multimodal factual dataset, it reveals that existing VLMs face significant challenges with questions requiring both detailed perception and factual knowledge. Moreover, open-source VLMs still demonstrate a substantial performance gap compared to proprietary ones. To advance knowledge-intensive VLMs, we conduct fine-tuning experiments using our development set, which improves LLaVA-1.5 evaluation accuracy by up to 3.1%. We hope that AgMMU can serve both as an evaluation benchmark dedicated to agriculture and a development suite for incorporating knowledge-intensive expertise into general-purpose VLMs.
Figure 1: AgMMU is a multimodal knowledge-intensive dataset with expertise data from the agricultural domain, for which vision-language models (VLMs) must observe the details of images and provide precise and fact-based answers. Originally derived from real-world user-expert conversations, AgMMU consists of five knowledge types, including insect identification, species identification, disease categorization, management instruction, and symptom description questions. From the question-answering perspective, AgMMU features 5,460 multiple-choice questions (MCQs) like conventional vision-language benchmarks [53] and open-ended questions (OEQs) like SimpleQA [50]. We also carefully curate an agricultural knowledge base with 205,399 pieces of information for foundation model fine-tuning. We hope AgMMU can benefit both knowledge-intensive VLMs and the social good.
[…]
.tocmtchapter
1 Introduction
Recent progress in large language models (LLMs) and vision language models (VLMs) has demonstrated remarkable capabilities in general knowledge reasoning and understanding, as exemplified by great performance on a variety of multimodal reasoning benchmarks [53, 36, 20, 50, 34, 28]. However, existing benchmarks often emphasize general-domain tasks and may not fully capture the performance limitations of these models in more specialized, knowledge-intensive domains. Generating factually accurate answers has recently attracted attention in language models, such as SimpleQA [49] proposed by OpenAI, but it still calls for studies on multimodal models in knowledge-intensive domains.
Agriculture, a cornerstone application of scientific biological knowledge, presents a particularly challenging domain and has been extensively investigated in the vision and machine learning community since early years [16, 27, 38, 15, 18, 13]. Unlike common tasks in the general domain, agricultural problems often require both precise visual analysis (e.g., identifying specific diseases, pests, or plant conditions) and extensive domain knowledge (e.g., treatment protocols, growing conditions, and management practices). Visual-based data are exceptionally crucial for plant management, because many serious plant problems (e.g., pests, disease, nutrient stress, water stress) are recognized visually and often with advice from remote experts. The stakes are also high in this scenario: accurate and timely agricultural diagnostics can mean the difference between crop survival and failure, directly impacting food security and farmer livelihoods. Despite the critical nature of this domain, we lack comprehensive benchmarks to evaluate and improve the agricultural reasoning abilities of VLMs.
The challenge of creating such benchmarks is multifaceted. First, biological and agricultural tasks, where computer vision is often applied, are notably labor-intensive. They require expert-labeled data specific to the target task. Meanwhile, agricultural expertise is scarce and specialized, making it difficult to curate high-quality evaluation data  [8]. Second, real-world agricultural problems are inherently multimodal and require both visual and textual understanding. Third, we do not have a clear protocol for what defines a representative distribution of realistic agricultural questions to be evaluated. These challenges have left a significant gap in our understanding of how well current AI systems can handle real-world agricultural problems. At the same time, there is strong current interest in LLM-based question answering services for agriculture, with agricultural chatbots being released by several large and small companies [11, 5, 39, 6, 9] and research projects [2].
In this work, we introduce the AgMMU (Agricultural Multimodal Understanding and Reasoning Benchmark), the first comprehensive benchmark designed to evaluate the capabilities of multimodal foundation models on agricultural knowledge and visual reasoning. Our benchmark leverages 116,231 real-world conversations between 2013-2024 on the Internet, which offers one-to-one expert answers from Cooperative Extension/University staff and volunteers within participating Land-Grant institutions [14]. This long-form real-world question-answering data source ensures that our benchmark captures authentic challenges faced by farmers and gardeners, including complex visual symptoms, varied environmental conditions, and nuanced management decisions. We design a multi-step pipeline to construct our dataset, which is composed of conversation categorization, knowledge extraction, and evaluation question generation procedures. We leverage large multimodal foundation models [10, 1] to process the raw information. The final dataset consists of a wide range of plant-related questions and factual knowledge, including disease and insect identification, recognition of plant species, management instruction, and symptom description tasks. An illustration of our benchmark is shown in Figure 1.
Evaluations conducted on a diverse collection of vision language models have revealed that existing large-scale foundation models are lacking in knowledge-extensive scenarios, suffering mainly from not having adequate knowledge to use nuanced situational context to conclude the correct answer. (refer to Sec. 4 for more details). Moreover, with our large multimodal agricultural knowledge base for fine-tuning, open-sourced model is able to achieve significant performance gain, exceeding strong closed-source models in almost all question subdomains, which highlight the significance of our curated.
Our contributions include: 1) AgMMU: A carefully curated evaluation set of 5,460 multiple-choice questions (MCQs) and open-ended questions (OEQs, where no candidate choices are provided), each extracted from QA pairs with expert answers, designed to test both visual understanding and knowledge application in agricultural contexts, and 2) AgBase-200K: A comprehensive agricultural multimodal knowledge dataset of 205,399 expert-verified facts and knowledge pairs, extracted from non-overlapping conversations, suitable for improving VLM performance on agricultural tasks. 3) A systematic evaluation of the leading VLMs on our benchmark, revealing critical insights about their current capabilities, limitations in handling knowledge-intensive agricultural queries, and systematic error analysis. Our analysis reveals a general phenomenon of VLMs in producing factually correct answers, and our AgMMU provides a complete training and evaluation suite for investigating such problems.
2 Related Work
Multimodal Foundation Models and Benchmarks.   The field of multimodal AI has witnessed remarkable progress in recent years, marked by increasingly sophisticated foundation models and comprehensive evaluation benchmarks. CLIP [32] marks early breakthroughs in large-scale multimodal foundation models and demonstrates strong zero-shot visual recognition capabilities. The emergence of VLMs has led to a new generation of models like LLaVA [24] that combine visual perception with sophisticated reasoning capabilities. Other work [46, 35, 25] advances the field by incorporating more careful dataset curation, better vision-language alignment, and stronger visual backbones. More recent developments have pushed the boundaries of scale and capability. Models like GPT-4 [1], Claude 3 [40], and Gemini [41, 33] have demonstrated remarkable general-purpose visual reasoning abilities.
The evolution of multimodal models has been accompanied by increasingly sophisticated evaluation benchmarks [22]. Early benchmarks like VQAv2 [17] and GQA [20] focus on basic visual question-answering capabilities. ScienceQA [34] specifically targeted scientific reasoning with visual components, while Eyes-Wide-Shut [47] evaluates the model’s ability to avoid hallucination. Recent benchmarks have emphasized real-world applications and complex reasoning. RealWorldQA [52] tests the ability of the models to handle practical, everyday visual queries. The MMMU [53] benchmark represents a comprehensive effort to evaluate multimodal understanding across multiple domains and task types. Unlike these methods, AgMMU presents the first comprehensive and challenging vision-language benchmark targeting agriculture, a domain that faces severe data scarcity and lacks domain-experts for large-scale dataset curation.
Computer Vision in Agriculture and Biology.   Computer vision has been extensively applied in the biological and agricultural domains [48, 54, 31, 3, 55], with significant datasets and benchmarks driving progress in species identification and disease classification [37, 12, 26, 51]. The iNaturalist dataset [18] marked a milestone by providing millions of species observations from citizen scientists, enabling the creation of a dataset and a deep neural network model (InsectNet) for the identification of pests [7, 13]. This was followed by more specialized collections like BioScan-1M/5M [15] which focused on microscopic biological images, and TreeOfLife-10M [38], which expands the data in a comprehensive phylogenetic framework. These datasets have facilitated numerous advances in automated species identification, disease detection, and biological image analysis. However, they cannot support complex real-world agriculture and gardening question answering.
Datasets Type Multimodal Training Expert Factuality
iNat21 [18] CLS -
TreeOfLife [38] CLS -
SimpleQA [50] OEQ
MMMU [53] MCQ
Our dataset MCQ+OEQ
Table 1: A comparison between AgMMU and some representative agricultural and general-domain datasets. Our proposed AgMMU is the first agricultural-domain multimodal question-answering datasets with evaluation and training sets. Moreover, our emphasis on expert domain and OEQs, instead of MCQs, uniquely challenges the factual knowledge precision of VLMs. (OEQ: open-ended question, MCQ: multi-choice question.)
3 The AgMMU Benchmark
Figure 2: Starting from raw user-expert conversations, we design a three-step data curation pipeline with carefully designed prompts and filtering. (1) Step 1. We employ LLaMA-70B to categorize the conversation and filter out the samples that do not fall under our selected agriculture sub-domains. (2) Step 2. A larger LLaMA-405B model extracts key pieces of agricultural knowledge from the long-form user-expert conversation. (3) Step 3. These facts either go to the evaluation set or knowledge base. For evaluation, the multiple-choice questions (MCQs) and open-ended questions (OEQs) are both formatted by GPT4o assisted with human quality verification.
AgMMU, short for “Agricultural Multimodal Understanding”, is a specially curated benchmark aimed at multimodal understanding in the agricultural domain. In addition to the importance of agriculture-related research, our AgMMU is also a novel benchmark for knowledge intensive multimodal understanding of general vision-language models, since agriculture-related questions generally require precise comprehension of image details (e.g., identification of pests, diseases, or weeds) and accurate memorization of facts (e.g., providing management suggestions).
To support studies on agricultural and knowledge-intensive VLMs, our AgMMU constructs a benchmark that contains both an agricultural knowledge base for training and an evaluation set with multiple-choice and open-ended questions. All data for AgMMU are collected from the Internet between 2013 and 2024, where users with agricultural questions communicate with agricultural experts from cooperative extension and university staff within land grant institutions [14]. These questions cover a wide range of plant knowledge, including weed management, pest control, general growth advice, generic plant identification, environmental stress management, and nutrient deficiency management. Examples are demonstrated in Figure 1.
Our AgMMU comprises two sets. (1) An evaluation set aiming to rigorously assess the abilities of existing models. It consists of 5,460 multiple-choice questions (MCQs) with curated choices, similar to the format of MMMU [53], and open-ended questions (OEQs) focusing on knowledge in short phrases, similar to SimpleQA [50]. (2) A large-scale knowledge base AgBase-200K supporting the training and development of models. It contains 205,399 pieces of facts extracted from 108,231 high-quality user-expert conversations. We also demonstrate that this knowledge base can be used to fine-tune existing VLMs.
Compared with related datasets in Table 1, our AgMMU has multi-facet advantages for developing VLMs. Regarding the biological and agricultural benchmarks of iNaturalist [18] and TreeOfLife-BioCLIP [38], AgMMU supports question-answering evaluation, which is crucial for a real-world VLM. As for general vision-language understanding benchmarks like MMMU [53], our dataset shares a similar emphasis as SimpleQA [50] on expert knowledge and factuality for VLMs. The large-scale training set AgBase-200K also distinguishes our dataset from common evaluation benchmarks in enabling expert-domain fine-tuning. As a combination of the above properties, AgMMU can uniquely benefit the development of both agricultural expert VLMs and the factual precision of general VLMs.
3.1 Dataset Curation
Curation Pipeline.   Each real-world conversation we collect contains a series of diverse facts. To convert a conversation into atomic QA pairs suitable for the evaluation of VLMs, we design a three-step pipeline as in Figure 2. (1) Categorization. We first tag user-expert QAs with a horticulture or agriculture domain and a primary plant organism of interest. (2) Knowledge Extraction. Based on the domain tag, we further decompose each expert-user conversation into its most important pieces of agricultural knowledge. (3) QA Generation. This step is for the evaluation set. We convert extracted agricultural facts into multiple-choice questions (MCQs) or open-ended questions (OEQs).
Step 1: Categorization.   Real-world agricultural questions come in a great variety with an open-ended nature. To extract clear agricultural knowledge, we must first account for the inherent complexity of these questions. Therefore, our first step is to tag each question with an agriculture sub-domain. The seven categories we decide on are [[[[disease advice”, “weed management”, “pests control”, “growing advice”, “environmental stress”, “nutrient deficiency”, “generic identification”, and other]]]]. We chose these categories based on the main categories designed by gardening and agriculture sites [42, 45, 44, 29, 30]. In this step, we filter out conversations in the “growing advice” and “other” sub-domains, as these are typically not image-related questions. Notably, the type of knowledge in user questions and expert answers often depends significantly on the question’s sub-domain, which we use to guide the next step, where prompts are selected according to the specific agricultural sub-domain.
Step 2: Knowledge Extraction.   In step 2, we extract agricultural knowledge available for training and evaluating VLMs. Across all questions, the user-expert conversation generally comprises a combination of identification (e.g., disease, or plant species) and management instructions. For example, in Figure 2, the user provides species identification information and symptom description, while the expert recognizes the diseases and offers management suggestions. To systematically extract knowledge from these conversations, we organize agricultural knowledge into categories of “species identification”, “disease identification”, “symptom/visual description”, “management instructions”, “insect/pest identification.”
To increase the reliability of information extraction, we leverage the agriculture sub-domain predicted in the previous step and design different prompt templates for each sub-domain. Our prompt leverages the following techniques for information extraction at higher qualities. (1) Our pipeline provides manually annotated in-context examples, significantly improving knowledge extraction quality and avoiding hallucination. (2) Our prompt explicitly verifies whether the expert has provided facts relevant to the target category, enabling the dynamic inclusion of available facts. We discover that LLMs [10, 1] can reliably complete the information extraction step. More details of our prompts and observations are available in the supplementary materials.
Step 3: QA Generation.   For rigorous evaluation, AgMMU adopts multiple-choice questions (MCQs) for all the knowledge types and open-ended questions (OEQs) for facts that are in the form of short phrases, inspired by MMMU [53] and SimpleQA [50], respectively. Note that only a small balanced set of knowledge is used for evaluation, while the rest goes to the knowledge base AgBase-200K. We use GPT-4o [1] to format agricultural facts extracted from previous steps into QA pairs.
We begin by carefully selecting a subset of samples for evaluation through a rigorous filtering and balancing process. (1) Filtering is done by removing the conversations without management information or issue identification, as we empirically find that the absence of such information indicates expert uncertainty. We also filter out questions with uncertain terms, such as “unclear” in the extracted information. (2) We remove questions with management instruction terms that direct users to an external source rather than giving them an actual answer, typically containing the words “lab,” “arborist,” etc. (3) To balance between different agriculture sub-domains (step 1) and knowledge types (step 2), we prioritize an even distribution across the types of questions, while maximizing the diversity of categories, as shown in Figure 3. This process extracts 8000 facts with high-quality and balanced distribution for evaluation.
Figure 3: Statistics of AgMMU. (1) The agricultural sub-domain distribution of our raw dataset, after the categorization step, as explained in Figure 2. (2) AgMMU, after the knowledge extraction and evaluation curation steps, serves as a balanced subset of raw dataset with proportional representation across knowledge types.
Then we employ GPT [1] to convert these facts into MCQs and OEQs. Depending on the knowledge types of the fact, we give the LLM a unique prompt to construct a corresponding question. We prompt it to generate three wrong answers and use the extracted fact as the ground truth to form one correct answer. To ensure fairness of the question, we extract “background information” with LLaMA-405B from the original AskExtension question which contains relevant information that the user provided. We also include the month,and location information in the question. This process gives us 5,460 unique MCQs and OEQs paired with 12,481 unique images.
Ethical Compliance.   We use automatic procedures assisted with human verification to erase the information that could leak the name, gender, address, or any other personal information for both users and experts. We will only distribute the processed AgMMU instead of the source data from the Internet to protect the privacy of the users and experts. To the best of our verification, we have not found any human faces in the benchmark images. Since we only keep the questions related to agriculture, our benchmark does not contain gender, race, or other types of societal biases.
3.2 Specialty of AgMMU
Distribution and Coverage.   In Figure 3, we show the distribution of agriculture sub-domains and knowledge types in AgMMU. We have two key observations: (1) The original dataset we obtained from the Internet is severely skewed in sub-domains: Sub-domains like nutrient deficiency and generic identification domains have significantly less samples than dominating domains such as disease advice or growing advice. This imbalance is inevitably carried over to knowledge types after the knowledge extraction stage. (2) However, in AgMMU, we demonstrates a much more balanced distribution with our dataset curation pipeline. This is exceptionally important for our evaluation set, where we aim for proposing an all-encompassing benchmark that examines different faces of models without inductive bias.
Realistic Images.   Our AgMMU also features significant challenges in the quality and number of images. The images in our benchmark are generally uploaded by users in various quality, resolution, and aspect ratios. This differs from the photography-level images in datasets like iNaturalist [18]. As shown in Figure 1, AgMMU realistically reflect the multimodal everyday scenarios encountered by gardeners and farmers. Our benchmark also exhibits the necessity of multi-image understanding challenge, since the user uploaded several images and the experts rely on a joint reasoning of them to make the final conclusion.
AgMMU-OEQs $\uparrow$ AgMMU-MCQs $\uparrow$
Disease/Issue Id.
Insect/Pest Id.
Species Recog.
Management Instr.
Symptom/Visual Descr.
Average
Disease/Issue Id.
Insect/Pest Id.
Species Recog.
Management Instr.
Symptom/Visual Descr.
Average
Proprietary Models
GPT-4o [1] 21.02 26.13 19.79 10.37 11.65 17.79 80.51 84.46 86.23 90.42 84.64 85.25
Gemini 1.5 Pro [33] 25.17 29.77 24.88 14.15 15.91 21.98 76.25 79.92 81.07 88.08 76.94 80.42
Claude 3 Haiku [40] 10.51 18.09 20.34 1.63 15.73 13.26 62.14 68.54 49.89 81.76 52.03 62.00
SOTA Open-sourced Models
Cambrian-8B [46] 6.78 17.29 18.55 1.96 5.00 9.92 63.66 60.84 57.49 73.28 76.38 65.81
InternVL2-8B [43] 4.20 15.61 8.13 1.79 6.64 7.27 50.95 58.14 58.34 73.13 61.31 60.17
LLaMA-3.2 [10] 11.56 23.96 21.25 4.02 18.77 15.91 65.07 69.29 75.21 78.50 78.08 73.32
LLaVA-NeXT-8B [25] 4.81 15.98 20.56 0.73 1.86 8.79 62.18 67.18 60.03 75.31 71.19 66.71
LLaVA-OneVision-8B [21] 5.97 16.02 15.39 2.59 8.90 9.77 69.07 70.93 67.62 81.87 75.02 72.53
LLaVA-1.5-7B [24] 3.98 14.64 23.07 0.97 9.45 10.42 62.66 61.24 57.18 70.64 71.32 64.16
LLaVA-1.5-13B [24] 5.19 15.70 24.21 2.58 10.72 11.68 66.57 58.73 60.41 78.29 72.26 66.73
Qwen-VL-7B [4] 4.97 17.94 22.34 1.89 8.24 11.08 56.24 58.31 60.65 69.64 67.70 62.34
VILA1.5-13B [23] 10.32 21.12 23.58 2.98 11.02 13.80 60.18 62.76 60.71 70.78 61.63 63.00
Table 2: Performance of VLMs on our AgMMU OEQs and MCQs. Our evaluation set poses great challenges to existing large VLMs, including closed-source and open-sourced models.
4 Experiments
We perform comprehensive evaluations on diverse vision-language models (VLMs), including both proprietary and open-sourced models. Our evaluation primarily focuses on zero-shot performance, utilizing either publicly available APIs or author-provided checkpoints to reflect the inherent capabilities of each model without task-specific training. Furthermore, we present fine-tuning results on LLaVA-1.5 to demonstrate the importance of our knowledge base AgBase-200K for optimizing performance on agriculture tasks. All experiments are conducted using NVIDIA A6000 GPUs, and further details regarding model configurations are available in the supplementary material.
4.1 Baselines and Metrics
Baselines.   We include a broad range of state-of-the-art VLMs to ensure robust comparisons. For open-sourced models, we prioritize architectures with comparable parameter numbers for a fair and meaningful comparison. Moreover, we include the most recent, largest, and highest-performing checkpoints up to the date of our evaluation questions, including LLaMA-3.2 [10], LLaVA-1.5 [24], LLaVA-NeXT [25], LLaVA-OneVision [21], Cambrian-1 [46], InternVL-2 [43], Qwen-VL [4], and VILA [23], as well as proprietary models such as GPT-4o [1], Claude-3.5 [40], and Gemini1.5-Pro [41]. Detailed configurations and model settings are in the supplementary material.
Evaluation.   We evaluate the models on multiple-choice questions (MCQs) and open-ended questions (OEQs). To ensure a robust and fair evaluation, we have developed a systematic approach to minimize potential variations in model responses due to intermediate generations, capitalization differences, or formatting inconsistencies.
For MCQs, we report accuracy as the primary evaluation metric for MCQs within the benchmark, reflecting the models’ ability to perform choice tasks relevant to agricultural scenarios. We score model response by conducting basic string pre-processing and matching the model letter prediction with the true letter associated with the ground truth.
For OEQs, we implement the LLM-as-judge methodology using Qwen-7B to grade the answers. The grading is divided into two categories: (1) few-word responses and (2) multi-statement responses. Few-word responses consist of the categories pest identification, disease identification, and species identification. These responses consist of two to three words and are evaluated directly against the correct answer. We use a structured scoring rubric following SimpleQA [50] and assign the grades of “correct”, “incorrect”, “partially correct”, and “not attempted”. We use the harmonic mean to calculate the final grade. Multi-statement responses, on the other hand, consist of the categories management instructions, and symptom descriptions. These categories cannot simply be evaluated directly against the correct answer because they usually consist of multiple unique and standalone statements. On average, the type of management instruction of responses has 2.73 correct unique statements and symptom/visual description type responses has 2.74 statements. Since we find the complexity of separating answers into unique statements and grading them individually to be too complex for a single inference of our LLM judge, we split the process into two steps. (1) In the first step, we ask the LLM judge to split the answer into standalone statements. And (2) in the next step, we have the LLM judge compare every statement in the correct answer with every statement in the predicted response using a similar grading rubric, which assigns the grades “correct”, “incorrect”, “partially correct”, and “irrelevant”.
The highest grade assigned to a predicted statement is saved as its score. For each question, we normalize the correct and partially correct predicted statements with the total correct statements. Then we use the harmonic mean to calculate the final grade [50].
4.2 Zero-shot Evaluation Results
We present a comprehensive zero-shot comparison of various VLMs in Table 2. More visualizations are presented in the supplementary material.
Very Challenging for Existing Models  : The AgMMU benchmark proves to be notably challenging across all evaluated models, and even the most advanced systems achieve moderate performance levels. This demonstrates that AgMMU reflects the inherent complexity of visual understanding and reasoning tasks in agriculture, where models must simultaneously process intricate visual details and apply domain-specific knowledge.
Performance Across Different Question Types  : For MCQs, the most challenging knowledge type categories are consistently those involving insect and disease identification, or species recognition. Management instruction and visual description questions present slightly smaller challenges, with most VLMs’ scores often 5-10% above the more complex recognition tasks. For OEQs, the management instruction and visual description types of questions pose more challenges, due to their long-response nature and the need for more reasoning. However, there is still significant variation between models as to what tasks are the most difficult, suggesting that specialized agricultural knowledge and visual understanding capabilities are not uniformly distributed across model architectures.
Figure 4: Error Analysis. The most common errors made by VLMs are knowledge errors. CoT represents samples that are originally wrong due to false reasoning, but corrected with chain-of-thought prompting. IPI, DII, SVD, SI, and MI are short for five question types as explain in Sec 3.1.
Open-ended Question Performance  : The performance difference between OEQs and MCQs is substantial, with the performance of the OEQ trailing 40 to 50% behind the MCQ results. These results suggest that free-response tasks, which do not allow the elimination process with candidate answers, require a much more nuanced understanding and generation of information and remain a considerable hurdle for open-sourced and closed-source models.
Between Closed- and Open-sourced Models  : We observe a distinct performance gap between closed-source and open-source models, where proprietary models generally demonstrate superior performance. Among all models, Gemini 1.5 Pro and GPT-4o lead in all tasks. For open-source models, LLaMA-3.2 is the overall best performer across all subdomains except management instructions where Cambrian-8B is the best. On average, open-sourced models trail behind their closed-source counterparts by more than 10-20%.
Figure 5: Error Demonstration. We show two common evaluation errors, including the lack of knowledge to answer the question (left), and the wrong perception of the image (right).
4.3 Error Analysis
We analyze the types of error that occurred during the VLM evaluation. We select 150 wrong evaluation samples and ask the VLM to provide the chain-of-thought reasoning before providing the final answer, and judge the error types with human evaluators. Figure 4 demonstrates quantitative results, and Figure 5 shows two common error types.
Knowledge Error (56.5%).   A knowledge error occurs when the VLM does not have the knowledge necessary to reach the correct answer given the context. In these cases, the model may identify relevant visual features and reinforce important context but cannot conclude the correct answer. For example, in management questions, knowledge errors often lead to vague or non-specific responses that fail to engage with the particular nuances of the given situation.
Figure 6: Open-ended questions (OEQ) examples demonstrating the effectiveness of our fine-tuned LLaVA model, which accurately identifies issues that both GPT-4o and the original LLaVA fail to recognize in zero-shot scenarios. In diverse agricultural categories: disease, species, and pest identification, our fine-tuned model exhibits enhanced diagnostic accuracy.
Figure 7: Finetuning on our knowledge base AgBase-200K significantly benefits the agricultural knowledge and factuality of LLaVA, indicating the potential of our training set for developing knowledge-intensive VLMs.
Perceptual Error (18.5%).   A perceptual error occurs when the VLM fails to recognize a primary visual characteristic or gives a directly incorrect description of a visual characteristic. This category highlights the incapability in visual understanding of VLMs.
Chain-of-Thought (22.8%).   The chain-of-thought (CoT) category includes instances where the VLM initially provides an incorrect response, but corrects itself when asked to explain its reasoning. The amount of samples in this category emphasizes the importance of reasoning procedures in obtaining the correct answers for many questions on the AgMMU benchmark.
Reject to Answer (2.2%).   In rare situations, we find that the VLM refuses to answer the question.
4.4 Finetuning Evaluation Results
[…]
Analysis.   Figure 6 and 7 demonstrate the effectiveness of fine-tuning a VLM using our AgBase-200K on more challenging open-ended questions. Figure 7 shows that the fine-tuning with our knowledge base brings a significant and consistent improvement across all question domains, leading to an overall 4.7% average improvement. This performance boost highlights the effectiveness of our large-scale knowledge base in improving VLMs in connecting facts with images. Compared to Table 1, having such a training set is an important benefit of our dataset. With the large-scale database, AgMMU is the first robust dataset for such domain-specific fine-tuning. The substantial performance gains observed emphasize the critical role of the dataset in improving the understanding of the model of complex domain-specific visual input and reasoning tasks.
5 Conclusion
In this work, we introduced AgMMU, the first benchmark for evaluating vision-language models (VLMs) in agriculture, a field that requires precise visual interpretation and expert knowledge. Our dataset spans core agricultural tasks, including symptom recognition, species and pest identification, and management instructions, and is created from more than 116,000 real-world expert-user interactions. Through a three-step curation framework, we build a high-quality AgMMU evaluation set, demonstrating a great challenge for current VLMs. To support model development, we introduce AgBase-200K, a training set of 205,399 knowledge entries aimed at improving model accuracy. We hope that the combination of AgMMU and AgBase can support the community in evaluating and developing stronger knowledge-intensive VLMs.
Acknowledgments.
We thank Kastan Day and Rohan Marwaha for the fruitful discussion. We thank the AskExtension team for helping us with the data curation. This work was supported in part by the AIFARMS National AI Institute, USDA-NIFA Award 2020-67021-32799, NSF Grant 2106825, the University of Illinois Discovery Partners Institute, the Center for Digital Agriculture at the University of Illinois, the Amazon-Illinois Center on AI for Interactive Conversational Experiences, and the IBM-Illinois Discovery Accelerator Institute. This work used computational resources, including the NCSA Delta and DeltaAI supercomputers through allocations CIS230012 and CIS240133 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, as well as the TACC Frontera supercomputer, Amazon Web Services (AWS), and OpenAI API through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
Supplementary Material
[…]
A Evaluated Large Vision Language Models
Our evaluation and analysis are conducted mainly on the group of models listed in Table 2 in the main paper. We have chosen models such that they cover most of the popular and best-performing methods used by recent multimodal understanding and reasoning work. In this part, we discuss all the models we have used in our experiments and explain their evaluation details, the public checkpoints we have chosen, and display the prompts we used to adapt the model to our datasets.
During evaluation, we chose to follow the standard prompt provided by the authors whenever possible for multiple-choice and short-answer questions. When the prompt is not provided for the model, we select a custom prompt that is created through several iterations of prompt engineering to select the one that produces the most effective results. For all open-source models, we only chose to only include the first associated image to each question even though some models support multi-image inference. The image is always included as the prefix. This way, we remain consistent and can reproduce the results.
Proprietary Models  . We used three proprietary models in our evaluation: GPT-4o [1], Gemini 1.5 Pro [33], and Claude 3 Haiku [40]. Below we note the model version and access date used for evaluation.
  • GPT-4o: November 1-13, 2024.
  • Gemini 1.5 Pro: November 1-13, 2024.
  • Claude 3 Haiku: November 13-14, 2024.
BioCLIP   [38]. BioCLIP is trained on Tree-Of-Life 10M, a large biological species dataset that allows it to learn hierarchical species names and relationships. By learning the hierarchical species relationship, it has greater generalization to species that were not in the training data, but closely related. We use the official implementation and checkpoint [^1] https://huggingface.co/imageomics/bioclip with a CLIP ViT-B/16 base model in our evaluation. During evaluation, we follow the evaluation methods of the original paper. Additionally, we only evaluated this model on Disease/Issue Identification, Insect/Pest Identification, and Species Recognition categories due to its major focus of species classification.
Cambrian-1   [46]. Cambrian-1 is a recent state-of-the-art model that excels at visual-centric tasks. This model explores combinations of vision encoders, text and image integration techniques, and instruction tuning strategies. We use the official implementation and checkpoint [^2] https://github.com/cambrian-mllm/cambrian with a LLaMA3-8B-Instruct LLM backbone model in our evaluation.
InternVL2   [43]. InternVL scales up the vision foundation model while aligning it with the backbone LLM, and is trained on web-scale image-text data to achieve strong performance across a variety of vision-centric tasks. We use the official implementation and checkpoint [^3] https://huggingface.co/OpenGVLab/InternVL2-8B with the InternViT-300M-448px vision backbone and Internlm2.5-7B-chat language backbone in our evaluation.
LLaMA-3.2   [10]. LLaMA-3.2 is the first collection of multimodal large language model from the LLaMA family that was previously text-only. The integration of vision involves utilizing cross-attention layers and a pre-trained vision encoder that feeds directly into the text-processor. The model follows a commonly used training recipe that includes pretraining on noisy image-text pairs and then high-quality knowledge enhanced pairs. Notably, the language-model parameters were frozen during the training of alignment of image and text to retain strong text-only capabilities. We use the official implementation and checkpoint [^4] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision that uses a LLaMA-3.1 text-only language backbone in our evaluation. When evaluating the model, we choose to use a custom prompt since no standard prompt is provided.
LLaVA-NeXT   [25]. LLaVA-NeXT expands on LLaVA by using the same instruction tuning method to give the model the ability to process and reason about multi-images, multi-grames, and multi-views. We use the official implementation and checkpoint [^5] https://huggingface.co/llava-hf/llama3-llava-next-8b-hf with LLaMA-3-8B Instruct as the language backbone in our evaluation.
LLaVA-OneVision   [21]. LLaVA-OneVision builds on LLaVA-NeXT with the capability to analyze single images, multi-images, and video scenarios. Most impressively, it allows for video understanding through task transfer from images but this is not really explored in our evaluations. We use the official implementation and checkpoint [^6] https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf that uses a base architecture consisting of SigLIP-SO400M-Patch14-384 and Qwen2-7b in our evaluation.
LLaVA-1.5-7B / LLaVA-1.5-13B   [24]. LLaVA introduces the idea of instruction tuning a multimodal model with GPT-4 generated instruction-following data for associated images. This gives it the ability to achieve impressive abilities to act as an instruction-following general agent. We use the official implementation and checkpoint [^7] https://huggingface.co/llava-hf/llava-1.5-7b-hf [^8] https://huggingface.co/llava-hf/llava-1.5-13b-hf with a CLIP ViT-L/14 vision backbone and Vicuna1.5-7B / Vicuna1.5-13B in our evaluation.
Qwen-VL-7B   [4]. Qwen-VL is a large vision language model that has the ability to perform various vision-language tasks including image captioning, visual grounding and more, not only limited to question answering. This model is multi-lingual in Chinese and English and was pre-trained using an interleaved image-text technique. We use the official implementation and checkpoint [^9] https://huggingface.co/Qwen/Qwen-VL that uses Qwen-7B as the language backbone and CLIP ViT bigG/14 as the vision encoder in our evaluation.
VILA1.5-13B   [23]. VILA is trained using an enhanced pre-training method that involves interleaved visual language data. Additionally, during the supervised fine-tuning stage, the data includes text-only instruction data to help the model retain strong text-only capabilities. We use the official implementation and checkpoint [^10] https://github.com/NVlabs/VILA with a LLaMA3-8B LLM backbone and SigLIP-SO400M-Patch14-384 vision encoder in our evaluation.
B Dataset Curation Details
This section outlines the three-stage curation pipeline of AgMMU and describes the prompts designed for each question type and subdomain.
Figure A: Prompts for our three-stage data curation pipeline.
Figure B: Examples included in prompt during agriculture domain categorization (step 1)
Figure C: Examples included in prompt during knowledge extraction (step 2) based on agriculture subdomain type
Figure D: Prompts used for each subdomain during knowledge extraction (step 2).
Figure E: Prompts for our LLM as judge on short-answer OEQ.
Figure F: Prompts for the answer splitting step of our LLM as judge on multi-sentence (long-answer) OEQ.
Figure G: Prompts for the evaluation step of our LLM as judge on multi-sentence (long-answer) OEQ.
B.1 Stage 1: Question Categorization
In the first step, we employ the Llama-70B model [10] to categorize questions into predefined agriculture subdomains while identifying the primary living entity affected by the query. Our systematically crafted prompt (Figure A) guides the model to extract the most specific living entity mentioned, such as “apple tree” or “honeybee,” or to assign “none” when the entity is unclear or absent.
The subdomains include Disease, Weeds/Invasive Plant Management, Insect/Pest Control, Growing Advice, Environmental Stress, Nutrient Deficiency, Generic Identification, and Other. Each subdomain is succinctly defined within the prompt, with illustrative examples provided in Figure B to address ambiguous or edge-case scenarios. The prompt enforces a standardized output format, ensuring consistency with the inclusion of “ENTITY:” and “CATEGORY:” headers.
To enhance robustness, the prompt includes examples of complex or overlapping cases, ensuring accurate classification even for questions that span multiple subdomains or lack explicit details. By embedding these clarifications, the design supports reliable categorization across diverse agricultural contexts.
B.2 Stage 2: Information Extraction
In the second step, we design prompts to extract granular categories of information from agricultural questions. These categories are tailored to the specific subdomain identified in Step 1, ensuring that the extracted information is both relevant and actionable.
Weeds/Invasive Plants Management.   For the “weeds/invasive plants management” subdomain, the extraction focuses on: (1) Image Description, visual characteristics of the weed or invasive plant, (2) Management Instructions, actionable strategies for control, and (3) Miscellaneous Facts, contextual expert insights. The name of the weed itself is already extracted in Step 1. This ensures that the emphasis remains on descriptions, actionable measures, and expert knowledge.
Insects/Pests Control.   For this subdomain, the categories include: (1) Insect/Pest, identifying the pest in focus, (2) Image Description, visual traits of the pest or evidence of damage, (3) Management Instructions, guidance for mitigation, and (4) Miscellaneous Facts, contextual expert insights. The primary plant affected, if exists, is identified in Step 1, thus this step concentrates on pest-specific details, such as visual features or damage patterns, and the corresponding management strategies.
Nutrient Deficiency, Disease, Environmental Stress.   For these subdomains, we group them due to shared characteristics. The extracted categories are: (1) Disease/Issue Identification, specifying the underlying cause, (2) Symptom Description, observable signs such as discoloration or stunted growth, (3) Management Instructions, remediation or prevention strategies, and (4) Miscellaneous Facts, contextual expert insights. These subdomains are defined by their symptomatic presentation, the underlying conditions, and the need for targeted management interventions.
Growing Advice.   For this subdomain, the variability in question structure necessitates tailored extractions: (1) Succinct Question, a concise reformulation of the user query, (2) Succinct Answer, a precise response to the query, (3) Image Description, any relevant visual details, and (4) Miscellaneous Facts, contextual expert insights.
Importantly, besides distinguishing the extraction types, we also put different examples of pre-made knowledge extraction into the prompt, see Figure C. Prompts given to the model for each subdomain can be seen in Figure D.
The Miscellaneous Facts category is extracted across all subdomains but is not directly used in subsequent steps. Instead, it captures standalone expert information that can contextualize a user’s issue.
To optimize extraction accuracy, we distinguish between “Symptom Description” (used for nutrient deficiency, disease, and environmental stress) and “Image Description” (used for weeds/invasive plants and insects/pests). While these serve a similar purpose—capturing observable or visual details—they are unified under the term “Symptom/Visual Description” in subsequent steps to maintain consistency.
B.3 Stage 3: Question Generation
In the final step, the extracted agricultural facts are transformed into evaluative question-answer (QA) pairs, comprising multiple-choice questions (MCQs) and open-ended questions (OEQs) generated using GPT-4o. To enhance relevance, we exclude two knowledge types: (1) Growing Advice, as image content often lacks direct correlation with the user’s issue, and (2) Miscellaneous Facts, since these provide general context but do not directly relate to the user’s image. This refinement narrows the scope to five key knowledge types for downstream processing, including Disease/Issue Identification, Symptom/Visual Description, Management Instructions, Insect/Pest, and Species.
To ensure clarity and relevance, we employ a standardized prompt structure (see Figure A tailored to each knowledge type. While the core structure remains consistent, the phrasing explicitly references the specific knowledge type being addressed. This targeted design allows the prompts to focus on generating well-contextualized and relevant questions. For added precision, the prompts incorporate contextual details where applicable: (1) For species-related questions, only symptom/visual description information is referenced, ensuring the focus remains on observable traits, and (2) for symptom/visual-related questions, species information is used to provide context, helping to ground the questions in specific agricultural scenarios.
This contextualization ensures that the generated questions integrate both user-provided information and extracted context seamlessly. The result is a set of comprehensive and ”fair” evaluative questions, designed to effectively assess multimodal agricultural understanding and reasoning.
Figure H: GPT-4o accuracy with increasing MCQ options. Model performance on MCQs across different categories, comparing accuracy scores when varying the number of answer options (4, 5, and 6). We observe a 5-10% difference in accuracy across categories between the 4-option and 6-option configurations, with performance generally decreasing as the number of options increases.
C More Evaluation and Design Choices
LLM as Judge.   To perform evaluation on few-word and multi-statement OEQ responses, we implemented the LLM-as-judge methodology using Qwen-7B. Our prompts for few-word responses (Figure E) and multi-statement responses (Figure F, G) contain several in-context examples to guide the LLM to correctly categorize the answer as “correct”, “incorrect”, “partially correct”, and “irrelevant”.
Number of MCQ options.   To determine the optimal number of answer choices for our MCQs, we conducted an ablation study comparing GPT-4o’s accuracy when presented with four, five, and six options. For efficiency, we conducted this experiment on a subset of 821 questions, generating 5 wrong answers with GPT-4o. We randomly choose 3, and 4 wrong answers, for the four-choice and five-choice experiment, respectively, and take all choices for the six-choice experiment. While this limited subset may not capture the full variability of the dataset, it provides sufficient evidence to inform our design decisions. Due to the risk of process of elimination with MCQs, we believe that OEQs more accurately capture model performance.
The results, shown in Figure  H, indicate that accuracy decreases as the number of answer choices increases. Specifically, we observed a 5-10% reduction in accuracy between the 4-option and 6-option configurations. This trend suggests that the model might rely on a process of elimination when selecting an answer, making it more challenging to identify the correct response as the number of options increases. While the decrease in accuracy is not overly significant, we think it justifies our choice to use four options for MCQs.
D More Dataset Visualization
In Table I, we demonstrate more samples in AgMMU with questions and multiple choice answers.
Figure I: Additional visualization of samples in AgMMU. Ground truth selections of each question are highlighted in yellow.
E Limitations and Future Work
While our work makes significant contributions to agricultural benchmark development and MLLM evaluation through knowledge-intensive tasks, we acknowledge several limitations and identify promising directions for future research in this section.
Advanced Utilization of Training Data.   Although our curated dataset, AgBase-200K, has proven significant effectiveness for fine-tuning MLLMs [24] as shown in Section 4 and Figure 7, its potential extends beyond our current usage. As a comprehensive knowledge repository, the dataset presents opportunities for knowledge retrieval and augmented generation approaches. In particular, the development of vision-centric multimodal retrieval augmented generation (RAG) systems remains an under-explored yet promising direction. This alternative could enable more effective knowledge extraction and utilization from our dataset, potentially improving model performance on agricultural understanding tasks. We leave the exploration of these advanced techniques for future work.
Expanded Model Coverage and Evaluation Protocols.   While our current study encompasses several state-of-the-art and most commonly used MLLMs for zero-shot evaluation and fine-tuning analysis, we acknowledge that these represent only a subset of available multimodal architectures and methodologies. To enhance the robustness and generalizability of our findings, we plan to incorporate a broader spectrum of MLLMs. Additionally, we plan to conduct more extensive ablation studies and comparative analyses across different model scales and architectures. This comprehensive evaluation will provide deeper insights into the relative strengths and limitations of various approaches in agricultural understanding tasks.
F Societal Impact
We anticipate no negative societal impact of our work. Our dataset is ethically designed, respecting the privacy of Extension.org users by removing personal identifying information such as name, gender, and location. Additionally, we have verified to the best of our ability to ensure the removal of images that contain human faces. During dataset curation, we put in great effort to eliminate bias by creating a dataset representative of the original Extension.org questions as well as a balanced dataset across all question types.
Positive Impact:   We hope that the creation and release of this challenging vision-knowledge intensive dataset can support active research in this domain. Our comprehensive dataset is adapted from real-world conversations between users and experts, creating samples that are more representative of questions and images one may ask. This enables more accurate responses as demonstrated by our fine-tuning experiments. This dataset can be used to support the development of an agricultural vision language model that can provide users with instant assistance on various topics like insect/pest identification, disease categorization, and most importantly, management instructions. When properly used, these models have the potential to assist sustainability goals, prevent yield loss, and improve resource use.